A centrifuge accelerates from rest to 15,000 rpm in 240 s. What is its angular acceleration?
Answer #1

Change in angular velocity=final angular velocity-initial angular velocity=15000-0=15000 rpm=1570.796 rad/s.

Time=240 s.

So, angular acceleration=(change in angular velocity)/time=1570.796/240=6.54 rad/s2.